OpenStudy (nathaliama13):

An unknown number y is 12 more than an unknown number x. The number y is also x less than 17. The equations to find x and y are shown below. y = x + 12 y = −x + 17 Which of the following statements is a correct step to find x and y? Multiply the equations to eliminate y. Write the points where the graphs of the equations intersect the x-axis. Write the points where the graphs of the equations intersect the y-axis. Add the equations to eliminate x.

OpenStudy (nathaliama13):

@Michele_Laino Are you busy? can you please explain to me how I solve this?

OpenStudy (anonymous):

Add the equations to eliminate x.

OpenStudy (anonymous):

By eye ball, y = 29/2

OpenStudy (nathaliama13):

I understand now. Thanks so much @robtobey

OpenStudy (michele_laino):

If I add those equations, I get: \[\Large y + y = x + 12 - x + 17\] please simplify

OpenStudy (nathaliama13):

\[y ^{2} = 29/2 ?? \]

OpenStudy (michele_laino):

no, we have: \[\Large 2y = 29\]